# Spike No More: Stabilizing the Pre-training of Large Language Models

### v2

@@ 5.2 Varying Sequence Length
Figure 8: Loss curves of “Vanilla” with 350M parameters in validation data when we vary the input sequence length. We adjust the batch size to use the same number of tokens for the training of each model.
@@
Li et al. (2022) indicated that it is better to train with a short sequence at the early stage to stabilize the LLM pre-training. They justified their method based on the curriculum learning strategy. On the other hand, in this section, we provide the theoretical justification to their method in terms of the standard deviation of the self-attention layer.
As discussed in Section 3.2, in Vanilla, the smaller the standard deviation of each sub-layer is, the larger the gradient of its layer is. In addition, the standard deviation of the self-attention layer depends on the result of $\mathrm{softmax}$ function. We considered the maximum value case in Section 3.2. In contrast, in this section, we consider the minimum value case, i.e., the result of $\mathrm{softmax}$ is the uniform distribution. Let $L$ be the length of an input sequence. When the result of $\mathrm{softmax}$ is the uniform distribution, we multiply $1/L$ by each vector in $X$. Because $X$ is the pack of layer-normalized vectors, their variances become $(1/L)^{2}$ after the multiplication. Then, we compute the sum of the $L$ vectors whose variances are $(1/L)^{2}$. To obtain the variance after the summation, we multiply $L$ by the $(1/L)^{2}$ based on Equation (6). Formally, we obtain the minimum variance of $\mathrm{softmax}(Z)X$ with the following equation:
$\mathrm{min}(\mathrm{var}(\mathrm{softmax}(Z)X))$ $=\left(\frac{1}{L}\right)^{2}L$
$=\frac{1}{L}.$ (15)
The above minimum value is smaller as $L$ is larger. Thus, the larger $L$ is, the more probable Vanilla is to face the loss and gradient spikes. In other words, a small $L$, which means that we use a short sequence for training, mitigates spikes.
@@
We conducted experiments with varying $L$ from $128$ to $2048$. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 8 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., $1024$ and $2048$. Moreover, the spikes are more probable to occur at the early stage of the pre-training. We suppose that the result of $\mathrm{softmax}$ in the early stage is similar to the uniform distribution because input vectors at the early stage are similar to random vectors. In other words, we suppose that the result of $\mathrm{softmax}$ at the early stage is challenging to be a peaked distribution because input vectors are not trained sufficiently. Therefore, using a short sequence stabilizes the training at the early stage, as reported in Li et al. (2022).
5.3 RMSNorm
Some recent LLMs use the RMSNorm (Zhang & Sennrich, 2019) instead of the LN in their Transformers (Touvron et al., 2023). We discuss such an architecture in this section. Based on the proof in Appendix D in Xiong et al. (2020), we can obtain the order of the RMSNorm gradient norm with the following equation:
$\left\lVert\frac{\delta\mathrm{RMSNorm}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{\sqrt{d}}{||x||_{2}}\right).$ (16)
In short, the order of the RMSNorm gradient norm is equal to the one of LN described in Equation (12). Thus, we also face the exploding gradients issue discussed in this paper theoretically if we use RMSNorm.
Figures 9 and 10 show the loss values and gradient norms of the Vanilla configuration in Section 4 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 4888We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in a similar location to the ones of Vanilla. These empirical results also indicate that the RMSNorms have the same issue as LNs regarding the exploding gradients.
Figure 9: Loss curves of Vanilla and RMSNorm in validation data.
Figure 10: Gradient norms of Vanilla and RMSNorm.
@@ 6 Conclusion
This paper investigated why LLMs sometimes face loss spikes during their pre-training. Specifically, we focused on gradients of internal layers. We introduced two causes of the exploding gradients during LLM pre-training through theoretical analyses: the shortcut and LNs. We also revealed requirements to prevent the exploding gradients. To satisfy the requirements, we also proposed a method combining simple modifications to embeddings and scaled initialization, a widely used initialization method in LLM pre-training. Experiments indicated that the combination avoided loss spikes. Moreover, we demonstrated that the combined method can train with a relatively larger learning rate, and as a result, it can achieve better performance. We hope our theoretical analyses and proposed method avoid wasting valuable time and computational resources during the LLM constructions.

### v3

@@ 6.2 Varying Sequence Length
Li et al. (2022) indicated that it is better to train with a short sequence at the early stage to stabilize the LLM pre-training. They justified their method based on the curriculum learning strategy. On the other hand, in this section, we provide the theoretical justification to their method in terms of the standard deviation of the shortcut.
Figure 6: Loss curves of Vanilla with 350M parameters in validation data when we vary the input sequence length. We adjust the batch size to use the same number of tokens for the training of each model.
As described in Section 2.1, Transformers add the output of each sub-layer to the shortcut. Since the standard deviation of the self-attention layer tends to decrease with the length of an input sequence especially at the early stage [^11] See Appendix H for details., a long sequence tends to keep the standard deviation of the shortcut small. Therefore, the long sequence makes the pre-training of Vanilla more unstable.
@@
We conducted experiments with varying the length of the input sequence $L$ from $128$ to $2048$. To use the same number of tokens to update parameters, we adjusted the batch size. Figure 6 shows loss values of Vanilla with each $L$ configuration in the validation data. This figure shows that spikes occurred only in the large $L$, i.e., $1024$ and $2048$. Moreover, the spikes are more likely to occur at the early stage of the pre-training. Therefore, using a short sequence stabilizes the training at the early stage, as reported in Li et al. (2022).
7 Related Work
Stability
To stabilize trainings of Transformer-based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023). Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct an LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper. To stabilize the LLM pre-training, Le Scao et al. (2022) applied the layer normalization to the embedding layer. Zeng et al. (2023) used shrink embedding gradient technique (Ding et al., 2021). In this study, we theoretically proved that the layer normalization to the embedding layer controls the upper bound of the gradient norms of sub-layers when we use the widely used initialization method for LLMs (Nguyen & Salazar, 2019; Shoeybi et al., 2020), and thus, it stabilizes the pre-training. For the initialization methods, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their training. Zhang et al. (2019) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we use their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that the widely used initialization method (Shoeybi et al., 2020), which makes parameters small, is necessary to stabilize the LLM pre-training. Moreover, we proved that we can prevent the loss spike problem by making the standard deviation of embeddings close to $1$.
Efficiency
As shown in Table 3, our modification enables the pre-training with a relatively larger learning rate, and can achieve better performance. Thus, this study can be regarded as on the efficiency of LLM pre-training because our modification can construct a better LLM with a given budget. Strubell et al. (2019) and Schwartz et al. (2019) reported that recent neural methods require substantial computational costs, and thus, they argued that we have to explore a cost-efficient approach. Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computations in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
@@ 8 Conclusion
This paper explored why large language models (LLMs) sometimes experience loss spikes during pre-training. To provide evidence, we specifically focused on the gradients of sub-layers. We introduced an upper bound for the gradient norms through an analysis of the spectral norms of the Jacobian matrices for the sub-layers. We then theoretically identified two conditions for avoiding loss spikes: small sub-layers and large shortcut. To meet these conditions, we show that using the widely adopted initialization method for LLMs can make the sub-layer parameters small, and that embedding scaling or incorporating layer normalization into the embedding layer can make the standard deviation of each embedding close to $1$, resulting in large shortcut. Experimental results indicated that methods satisfying these conditions avoid loss spikes. Furthermore, these methods allow for training with a relatively larger learning rate, leading to improved performance. We hope our theoretical analyses and empirical findings will help avoid wasting valuable time and computational budgets during LLM construction.
Ethics Statement
To stabilize the LLM pre-training, this paper provides theoretical analyses om the spectral norms of the Jacobian matrices for sub-layers to estimate the upper bound of the gradient norm of $\mathcal{L}$. This paper focuses on only the stability of LLM pre-training, and thus, we have to address other issues of LLMs such as hallucinations to use the LLM in a real application.
Reproducibility Statement
We do not aim to propose a novel method in this paper, but we mainly focus on theoretical analyses on the spectral norms of the Jacobian matrices to find the factor to stabilize the pre-training of LLMs. We justify our theoretical analyses through experiments with various situations. To activate our modification, we add only several lines to a widely used implementation, i.e., Megatron-LM [^12] https://github.com/NVIDIA/Megatron-LM. Therefore, we believe that it is easy to reproduce our experimental results. However, because it is difficult to conclude that our provided conditions completely solve the instability during pre-training of LLMs, it is better to combine other techniques to stabilize the pre-training such as an auxiliary loss described by Chowdhery et al. (2022) to make the pre-training more stable in an actual pre-training situation.
@@ Appendix A Hyper-Parameters
Table 4 shows that hyper-parameters used in our experiments on LLMs. In addition to experiments described in Section 5, this table also indicates the hyper-parameters of the model with 13B parameters that we evaluated in Appendix C.